OpenStudy (anonymous):

i need some help with solving 3 integrals in terms of u. i will post them in a min

OpenStudy (anonymous):

\[\int\limits_{2}^{3}6x^2(x^3+1)^4 dx\] \[\int\limits_{\pi/12}^{\pi/2}x \cos(4x^2)dx\] \[\int\limits_{2}^{\sqrt{11}}x \sqrt{x^2+5} dx\]

OpenStudy (tkhunny):

Have you considered substituting u = "Whatever is inside the parentheses"?

OpenStudy (anonymous):

yes that is what i did but from there i dont know how to get it fully in terms of u

OpenStudy (tkhunny):

Demonstrate one. \(u = x^{3} + 1\) \(du = 3x^{2}\;dx\) Go!

OpenStudy (anonymous):

for the first one i would pull the 6 outside and multiply the inside by 3 and the outside by 1/3

OpenStudy (tkhunny):

?? Why add a factor of 3 when you already have one sitting there? \(u = x^3 + 1\) This is monotonic, increasing, so there is no confusion. \(2^{3} + 1 = 8 + 1 = 9\) \(3^{3} + 1 = 27 + 1 = 28\) Integrand: \(6x^{2}\left(x^{3}+1\right)\;dx = 2\left(x^{3}+1\right)\cdot 3x^{2}\;dx = 2u\;du\) \(\int_{9}^{28}2u\;du\)

OpenStudy (anonymous):

oh ok now i understand what i did wrong

OpenStudy (tkhunny):

Let's see the next one.